### v1

# Refining Ky Fan’s majorization relation with linear programming

Abstract
A separable version of Ky Fan’s majorization relation is proven for a sum of two operators that are each a tensor product of two positive semi-definite operators. In order to prove it, upper bounds are established for the relevant largest eigenvalue sums in terms of the optimal values of certain linear programs. The objective function of these linear programs is the dual of the direct sum of the spectra of the summands. The feasible sets are bounded polyhedra determined by positive numbers, called alignment terms, that quantify the overlaps between pairs of largest eigenvalue spaces of the summands. By appealing to geometric considerations, tight upper bounds are established on the alignment terms of tensor products of positive semi-definite operators. As an application, the spin alignment conjecture in quantum information theory is affirmatively resolved to the 2-letter level. Consequently, the coherent information of platypus channels is additive to the 2-letter level.
1 Introduction
Among the open problems in quantum information theory, the problem of regularized formulas sticks out like a sore thumb. Put briefly, known formulas for the optimal rates of asymptotically error-free information processing—a principal concern of the theory— are given by limits of sequences of optimizations over domains with exponentially growing dimensions. Often, little is known about the convergence rates of these sequences. A common surmise is that much is yet to be understood about information processing in quantum-mechanical contexts. To make progress, new analysis tools are necessary.
Let us consider the transmission of quantum information through a memoryless quantum channel $\mathcal{N}$ (a linear completely positive trace-preserving map). The quantum capacity of $\mathcal{N}$ is the largest rate of asymptotically error-free transmission of quantum information through $\mathcal{N}$, measured in qubits per channel use. Let $\mathcal{Q}(\mathcal{N})$ denote the quantum capacity of $\mathcal{N}$. We have one formula for $\mathcal{Q}(\mathcal{N})$, due to Lloyd [15], Shor [22] and Devetak [8], in terms of the coherent information of $\mathcal{N}$. Let $\mathcal{N}^{c}$ denote a complementary channel to $\mathcal{N}$ defined via a Stinespring dilation (see, for example, Ch. 5 of Ref. [23]). The output of $\mathcal{N}^{c}$ is conventionally called the environment of $\mathcal{N}$. The coherent information of $\mathcal{N}$ is given by
$I^{\text{(coh)}}(\mathcal{N}):=\max_{\rho}H(\mathcal{N}(\rho))-H(\mathcal{N}^{c}(\rho)),$ (1.1)
[…]
$\mathcal{Q}(\mathcal{N})=\lim_{n\rightarrow\infty}\frac{I^{\text{(coh)}}(\mathcal{N}^{\otimes n})}{n}.$ (1.2)
It follows from the additivity of the von Neumann entropy under tensor products that $I^{\text{(coh)}}(\mathcal{N}^{\otimes n})\geq nI^{\text{(coh)}}(\mathcal{N})$. If $I^{\text{(coh)}}(\mathcal{N}^{\otimes n})=nI^{\text{(coh)}}(\mathcal{N})$, then $\mathcal{N}$ is said to have additive coherent information to the $n$-letter level. If this is the case for all $n$, then $\mathcal{Q}(\mathcal{N})$ has the manifestly computable formula $I^{\text{(coh)}}(\mathcal{N})$. Unfortunately, this is generically not the case (see, for example, Sec. IV.B of Ref. [12]). A stark instance of this non-additivity is found in Ref. [7], where Cubitt et al. show that for every positive integer $m$, there exists a quantum channel $\mathcal{N}_{m}$ such that $I^{\text{(coh)}}(\mathcal{N}_{m}^{\otimes m})=0$ while $\mathcal{Q}(\mathcal{N}_{m})>0$.
Here we are concerned with a conjecture called the spin alignment conjecture. Introduced by Leditzky et al. in Ref. [14], it is a conjecture about the entropies of convex mixtures of quantum states that have a certain tensor product structure. It arose in the context of computing the quantum capacities of platypus channels. The particular form of these quantum channels is not relevant to us here. By a reduction in Sec. V of Ref. [14], each platypus channel $\mathcal{M}$ has an associated quantum state $M$ such that the coherent information of $\mathcal{M}$ is additive to the $n$-letter level if for all probability distributions $p$ on the subsets of $[n]:=\{1,2,\ldots,n\}$ and all tuples of pure (rank-$1$) quantum states $(\psi_{J})_{J\subseteq[n]}$, it holds that
$H(\sum_{J\subseteq[n]}p_{J}\psi_{J}\otimes M^{\otimes J^{c}})\geq H(\sum_{J\subseteq[n]}p_{J}(v_{M}v_{M}^{*})^{\otimes J}\otimes M^{\otimes J^{c}}),$ (1.3)
where $v_{M}$ is a unit eigenvector of $M$ corresponding to its largest eigenvalue. For each subset $J\subseteq[n]$, the expression $\psi_{J}\otimes M^{\otimes J^{c}}$ denotes a quantum state that is separable across the bipartition $J\mid J^{c}$. The factor $\psi_{J}$ denotes the quantum state of particles whose labels are in $J$ and the tensor power $M^{\otimes J^{c}}$ denotes the quantum state of particles whose label are in the complement $J^{c}$. The spin alignment conjecture is that the inequality above holds in general. Results concerning the interplay between platypus channels and other quantum channels are shown in Ref. [13], some of which are conditioned on this conjecture.
Described in words, the entropy of the convex mixture $\sum_{J\subseteq[n]}p_{J}\psi_{J}\otimes M^{\otimes J^{c}}$, which may be interpreted as the quantum state of the environment of $\mathcal{M}^{\otimes n}$, is conjectured to be smallest when $(\psi_{J})_{J\subseteq[n]}$ is chosen so that the particles have maximally aligned spins. The stronger conjecture that
$\sum_{J\subseteq[n]}p_{J}\psi_{J}\otimes M^{\otimes J^{c}}\preceq\sum_{J\subseteq[n]}p_{J}({v_{M}}{v_{M}}^{*})^{\otimes J}\otimes M^{\otimes J^{c}},$ (1.4)
where $\preceq$ is the majorization symbol (see Sec. 2 for a formal definition), was pursued in Ref. [1]. As of this writing, a complete resolution of either conjecture is yet to be found.
It is instructive to consider the $n=1$ case. For a self-adjoint (or Hermitian) operator $A$ on $\mathbb{C}^{d}$, $\lambda(A)\in\mathbb{R}^{d}$ denotes the tuple of eigenvalues of $A$ arranged so $\lambda_{1}(A)\geq\lambda_{2}(A)\geq\cdots\geq\lambda_{d}(A)$. The relation above immediately follows from Fan’s majorization relation [9] (also known as Ky Fan’s eigenvalue inequality [17]):
$\lambda(p_{\{1\}}\psi_{\{1\}}+p_{\varnothing}M)$ $\preceq\lambda(p_{\{1\}}\psi_{\{1\}})+\lambda(p_{\varnothing}M)$ (1.5)
$=\lambda(p_{\{1\}}v_{M}v_{M}^{*})+\lambda(p_{\varnothing}M)$ (1.6)
$=\lambda(p_{\{1\}}v_{M}v_{M}^{*}+p_{\varnothing}M).$ (1.7)
The last equality is crucial to this argument. More generally, for self-adjoint operators $A_{1}$ and $A_{2}$, the equality $\lambda(A_{1})+\lambda(A_{2})=\lambda(A_{1}+A_{2})$ holds if and only if there exist orthonormal vectors $w_{1},\ldots,w_{d}$ such that $A_{1}w_{i}=\lambda_{i}(A_{1})w_{i}$ and $A_{2}w_{i}=\lambda_{i}(A_{2})w_{i}$ for each $i\in[d]$ (see, for example, Ref. [20]). This invites the following definitions.
Definition 1.1.
Let $q_{1},\ldots,q_{d}$ denote elements of the standard orthonormal basis for $\mathbb{C}^{d}$. We define an alignment map on the space of self-adjoint operator on $\mathbb{C}^{d}$ by the action $X\mapsto X^{\downarrow}:=\sum_{i=1}^{d}\lambda_{i}(X)q_{i}q_{i}^{*}$.
Definition 1.2.
Given self-adjoint operators $X_{1},X_{2},\ldots,X_{m}$ on $\mathbb{C}^{d}$, we say that they are perfectly aligned if there exists a unitary operator $U$ on $\mathbb{C}^{d}$ such that $X_{i}^{\downarrow}=UX_{i}{U}^{*}$ for each $i\in[m]$.
By considering majorization as a preorder on the space of self-adjoint operators, Fan’s majorization relation may be written as follows:
$A_{1}+A_{2}\preceq A_{1}^{\downarrow}+A_{2}^{\downarrow}.$ (1.8)
This relation generalizes to conic combinations of self-adjoint operators.
The obstacle to extending the argument above to cases where $n>1$ is that, in general, there may not exist a pure quantum state tuple $(\psi_{J})_{J\subseteq[n]}$ such that the summands $p_{J}\psi_{J}\otimes M^{\otimes J^{c}}$ are perfectly aligned. However, at the conjectured optimal tuple $({v_{M}{v_{M}}^{*}}^{\otimes J})_{J\subseteq[n]}$, the quantum state of each particle is a convex combination of the perfectly aligned operators $M$ and $v_{M}v_{M}^{*}$. This suggests that local perfect alignment may yield minimum entropy when global perfect alignment is impossible.
The main contribution of this work is a formalization of this suggestion for a sum of two tensor products of two positive semi-definite operators.
Theorem 1.3.
Let $B_{1},B_{2}$ denote positive semi-definite operators on $\mathbb{C}^{d_{B}}$ and $C_{1},C_{2}$ denote positive semi-definite operators on $\mathbb{C}^{d_{C}}$. Then
$B_{1}\otimes C_{1}+B_{2}\otimes C_{2}\preceq B_{1}^{\downarrow}\otimes C_{1}^{\downarrow}+B_{2}^{\downarrow}\otimes C_{2}^{\downarrow}.$ (1.9)
We call this relation a separable Fan majorization relation. We use it to resolve the spin alignment conjecture to the $2$-letter level (Prop. 4.4). The crucial ingredient of our proof of this relation is an observation about the orthogonality of subspaces that are spanned by tensor product vectors (Prop. 4.2).
1.1 Structure of the paper
Sec. 2 contains notation and relevant mathematical background. Sec. 3 contains an estimation schema based on linear programming that refines Fan’s majorization relation. In Sec. 4, the schema is used to establish Eq. 1.9. The paper is concluded in Sec. 5. The appendix App. A contains a deferred part of a proof that the linear programming bounds are tight in the case of simultaneously diagonalizable summands.
2 Notation and background
We refer to a set of $k$ elements as a $k$-set. For $x\in\mathbb{R}^{d}$, $x^{\downarrow}$ denotes the vector in $\mathbb{R}^{d}$ whose coordinates are the coordinates of $x$ albeit ordered (weakly) decreasingly. For $k\in[d]$, the sum of the $k$ largest coordinates of $x$ is denoted by $s_{k}(x):=\sum_{i=1}^{k}x_{i}^{\downarrow}$. For a subset $T\subseteq[d]$, we write $\mathtt{1}_{T}\in\mathbb{R}^{d}$ to denote indicator function of $T$. That is, the vector that satisfies for each $i\in T,(\mathtt{1}_{T})_{i}=1$ and for each $i\notin T,(\mathtt{1}_{T})_{i}=0$. A vector $x\in\mathbb{R}^{d}$ is said to majorize a vector $y\in\mathbb{R}^{d}$ if
$s_{k}(x)\geq s_{k}(y)$ (2.1)
for each $k\in[d-1]$ and $s_{d}(x)=s_{d}(y)$. This statement is denoted by $y\preceq x$. Majorization is defined for self-adjoint operators spectrally (see Sec. 7 of Ref. [2]). If $A,A^{\prime}$ are self-adjoint operators on $\mathbb{C}^{d}$ and $\lambda(A^{\prime})\preceq\lambda(A)$, then $A$ is said to majorize $A^{\prime}$, and we denote that by $A^{\prime}\preceq A$. As shorthand, we write $s_{k}(A):=\sum_{i=1}^{k}\lambda_{i}(A)$. As a relation on a real vector space, majorization is reflexive and transitive, that is, a preorder. See Ref. [16] for a more comprehensive account of majorization theory.
The statement that a vector space $W$ is a subspace of a vector space $W^{\prime}$ is denoted by $W\leq W^{\prime}$. We denote the dimension of a subspace $W$ by $\dim(W)$ the orthogonal projection operator, or projector, onto $W$ by $P_{W}$.
Let $A$ be a self-adjoint operator on $\mathbb{C}^{d}$. For $k\in[d]$, Fan’s maximum principle, proven in Ref. [9], gives an extremal formula for the sum of the $k$ largest eigenvalues of $A$:
[…]
Fan’s majorization relation Eq. 1.8 is an immediate consequence of this formula. Let $\xi_{1}(A),\xi_{2}(A),\ldots,\xi_{d}(A)$ denote orthonormal vectors in $\mathbb{C}^{d}$ such that for each $i\in[d]$, $A\xi_{i}(A)=\lambda_{i}(A)\xi_{i}(A)$. These vectors and the corresponding eigenvalues give an orthonormal spectral decomposition of $A$. For brevity, we use $\Pi_{i}(A)$ to denote the rank-$1$ projector $\xi_{i}(A)\xi_{i}(A)^{*}$. For each $\ell\in[d]$, we denote the linear span of the first $\ell$ of these eigenvectors by
$V_{\downarrow\ell}(A):=\langle\{\xi_{i}(A)\}_{i=1}^{\ell}\rangle.$ (2.3)
We refer to $V_{\downarrow\ell}(A)$ as a largest $\ell$ eigenvalue space of $A$. We denote the projector onto this space by $P_{\downarrow\ell}(A)$. It is useful to express the data of such a spectral decomposition by a complete flag [10],
$V_{\bullet}(A)=(0\leq V_{\downarrow 1}(A)\leq V_{\downarrow 2}(A)\leq\cdots\leq V_{\downarrow d}(A)=\mathbb{C}^{d}).$ (2.4)
If $\lambda(A)$ contains repeated values, that is, $A$ is degenerate, then $V_{\bullet}(A)$ is not uniquely defined. In such a case, we assume a choice is fixed throughout the analysis.
A linear program is an optimization problem whose objective function is linear and whose feasible set is a polyhedron. We denote optimal points of linear programs by a bar over a variable symbol, for example $\bar{x}$. We write ‘s.t.’ as shorthand for the phrase ‘subject to’. See Ref. [21] for more details on the theory of linear programming.
Let $\mathcal{T}$ be a partially ordered set. A subset $T\subseteq\mathcal{T}$ is called a downward closed set if for all $x\in T$ and $y\in\mathcal{T}$, $y\leq x\implies y\in T.$ The dual notion, meaning under replacing the symbol $\leq$ with the symbol $\geq$, is called an upward closed set. The complement of a downward closed set is an upward closed set. Given a set $S\subseteq\mathcal{T}$, its downward closure $\downarrow S:=\{y\in\mathcal{T}\mid\exists\,x\in S,y\leq x\}$ is the smallest downward closed set that contains it. The upward closure $\uparrow S$ is defined dually. See J.B. Nation’s lecture notes on order theory [18].
3 Estimation schema
Let $A_{1},A_{2}$ be self-adjoint operators on $\mathbb{C}^{d}$ and fix $k\in[d]$. Our aim is to find upper bounds on $s_{k}(A_{1}+A_{2})$ that are generally tighter than $s_{k}(A_{1})+s_{k}(A_{2})$, the upper bound from Fan’s majorization relation Eq. 1.8.
For a $k$-dimensional subspace $W\leq\mathbb{C}^{d}$ and $i\in[d]$, we denote the overlaps
[…]
$=(\lambda(A_{1})\oplus\lambda(A_{2}))^{T}x_{W}.$
It can be verified by inspection that $x_{W}$ satisfies the linear constraints:
$x^{(1)}_{W,i},x^{(2)}_{W,i}\geq 0,\forall\,i\in[d],$ (3.3)
$x^{(1)}_{W,i},x^{(2)}_{W,i}\leq 1,\forall\,i\in[d],$
$\sum_{i=1}^{d}x^{(1)}_{W,i}=\sum_{i=1}^{d}x^{(2)}_{W,i}=k.$
We refer to these constraints as the basic constraints. The set of points in $\mathbb{R}^{2d}$ that satisfy the basic constraints is a bounded polyhedron that we denote by $\mathcal{P}_{0}^{(k)}$. The inequality $s_{k}(A_{1}+A_{2})\leq s_{k}(A_{1})+s_{k}(A_{2})$ can be proven in a rather fancy way using this polyhedron. Observe that
[…]
$=\max_{\dim(W)=k}(\lambda(A_{1})\oplus\lambda(A_{2}))^{T}x_{W}\leq(\lambda(A_{1})\oplus\lambda(A_{2}))^{T}\bar{x},$ (3.5)
where $\bar{x}\in\mathbb{R}^{2d}$ is an optimal point for the linear program
$\max\>(\lambda(A_{1})\oplus\lambda(A_{2}))^{T}x$ (3.6)
[…]
A routine calculation shows that $\mathtt{1}_{[k]}\oplus\mathtt{1}_{[k]}$ is an optimal point.
To obtain sharper estimates, we need to consider more constraints. To that end, we introduce alignment terms.
Definition 3.1.
For each pair $(\ell_{1},\ell_{2})\in[d]\times[d]$, we define the alignment term
[…]
$=s_{k}(P_{\downarrow\ell_{1}}(A_{1})+P_{\downarrow\ell_{2}}(A_{2})).$ (3.8)
These $d^{2}$ positive numbers are meant to capture the overlaps between the largest eigenvalue spaces of $A_{1}$ and the largest eigenvalue spaces of $A_{2}$. The flag pair argument $(V_{\bullet}(A_{1}),V_{\bullet}(A_{2}))$ is included to emphasize the dependence of alignment terms on the assumed spectral decompositions of $A_{1}$ and $A_{2}$. When it is clear from context, we omit it and write $\alpha_{(\ell_{1},\ell_{2})}^{(k)}$.
For any $k$-dimensional subspace $W\leq\mathbb{C}^{d}$, $x_{W}$ satisfies, by definition, the linear constraint
$(\mathtt{1}_{[\ell_{1}]}\oplus\mathtt{1}_{[\ell_{2}]})^{T}x_{W}\leq\alpha_{(\ell_{1},\ell_{2})}^{(k)}(V_{\bullet}(A_{1}),V_{\bullet}(A_{2})).$ (3.9)
We call this an alignment constraint. We denote the polyhedron of points in $\mathcal{P}_{0}^{(k)}$ that satisfy all alignment constraints by $\mathcal{P}_{1}^{(k)}(V_{\bullet}(A_{1}),V_{\bullet}(A_{2}))$ or simply $\mathcal{P}_{1}^{(k)}$.
Theorem 3.2.
Let $u_{k}(A_{1},A_{2})$ denote the optimal value of the linear program
$\max\>(\lambda(A_{1})\oplus\lambda(A_{2}))^{T}x$ (3.10)
[…]
Then $s_{k}(A_{1}+A_{2})\leq u_{k}(A_{1},A_{2})$.
Proof.
The bound follows from Fan’s maximum principle Eq. 2.2, the fact that $x_{W}\in\mathcal{P}_{1}^{(k)}(V_{\bullet}(A_{1})V_{\bullet}(A_{2}))$ for $k$-dimensional $W$, and the equalities in Eq. 3.2. ∎
Several remarks are in order.
  1. 1.
    Although $\alpha_{(\ell_{1},\ell_{2})}^{(k)}$ is given by a maximization over $k$-dimensional subspaces, we find it is more amenable to analysis than $s_{k}(A_{1}+A_{2})$. After all, the spectrum of a sum of two projectors is closely related to the relative position of their supports (see, for example, Ref. [11]). This is a crucial ingredient in our proof of Prop. 4.3.
  2. 2.
    If $\mathcal{P}\subseteq\mathbb{R}^{2d}$ is a polyhedron that satisfies $\mathcal{P}_{1}^{(k)}(V_{\bullet}(A_{1}),V_{\bullet}(A_{2}))\subseteq\mathcal{P}\subseteq\mathcal{P}_{0}^{(k)}$, then the optimal value of the linear program
    $\max\>(\lambda(A_{1})\oplus\lambda(A_{2}))^{T}x$ (3.11)
[…]
    is an upper bound for $s_{k}(A_{1}+A_{2})$ that is at most $s_{k}(A_{1})+s_{k}(A_{2})$. For example, $\mathcal{P}$ could be the polyhedron of points that satisfy the basic constraints and a subset of the alignment constraints. Or $\mathcal{P}$ could be equal to $\mathcal{P}_{1}^{(k)}(V_{\bullet}(\tilde{A_{1}}),V_{\bullet}(\tilde{A_{2}}))$ where $\tilde{A_{1}},\tilde{A_{2}}$ are self-adjoint operators on $\mathbb{C}^{d}$ such that
    $\alpha_{(\ell_{1},\ell_{2})}^{(k)}(V_{\bullet}(A_{1}),V_{\bullet}(A_{2}))\leq\alpha_{(\ell_{1},\ell_{2})}^{(k)}(V_{\bullet}(\tilde{A_{1}}),V_{\bullet}(\tilde{A_{2}}))$ (3.12)
    holds for each pair $(\ell_{1},\ell_{2})\in[d]\times[d]$.
  3. 3.
    The upper bound $u_{k}(A_{1},A_{2})$ is in fact independent of the choice of spectral decompositions for $A_{1}$ and $A_{2}$. To see this, let us suppose that for some $t,f\in[d]$, $\lambda_{t}(A_{1})=\lambda_{t+1}(A_{1})=\cdots=\lambda_{t+f}(A_{1})$. Let $U$ be a unitary operator on $\mathbb{C}^{d}$ that acts as the identity on $\langle\{\xi_{i}(A_{1})\}_{i=t}^{t+f}\rangle^{\perp}$ but is otherwise arbitrary. Let $V^{\prime}_{\bullet}(A_{1})$ denote the flag
    $(0\leq UV_{\downarrow 1}(A_{1})\leq UV_{\downarrow 2}(A_{1})\leq\cdots\leq UV_{\downarrow d}(A_{1})=\mathbb{C}^{d}).$ (3.13)
    For $X$ self-adjoint on $\mathbb{C}^{d}$, the change of basis effected by $U$ induces a linear action on the corresponding overlap tuple:
[…]
    For $y^{(1)}\in\mathbb{R}^{d}$, it follows from the fact that the restriction of $U$ to $\langle\{\xi_{i}(A_{1})\}_{i=t}^{t+f}\rangle$ is unitary that the sum $\sum_{i=t}^{t+f}y^{(1)}_{i}$ is invariant under this action. Now, suppose that $x^{(1)}\oplus x^{(2)}\in\mathcal{P}_{1}^{(k)}(V_{\bullet}(A_{1}),V_{\bullet}(A_{2}))$. Let $x^{\prime(1)}$ denote the vector with coordinates
    $(x^{\prime(1)}_{t},x^{\prime(1)}_{t+1},\ldots,x^{\prime(1)}_{t+f}):=(0,0,\ldots,\sum_{i=t}^{t+f}x^{(1)}_{i}),$ (3.15)
    and otherwise $x^{\prime(1)}_{i}:=x^{(1)}_{i}.$ It can be verified by inspection that $x^{\prime(1)}\oplus x^{(2)}\in\mathcal{P}_{1}^{(k)}(V^{\prime}_{\bullet}(A_{1}),V_{\bullet}(A_{2}))$ and that the objective value at $x^{\prime(1)}\oplus x^{(2)}$ equals the objective value at $x^{(1)}\oplus x^{(2)}$. These considerations extend to cases where there are other degeneracies in either $\lambda(A_{1})$ or $\lambda(A_{2})$.
3.1 Total unimodularity of the constraints
A useful observation about the alignment constraints is that vectors of the form $\mathtt{1}_{[\ell_{1}]}\oplus\mathtt{1}_{[\ell_{2}]}$ can be made to have the consecutive ones property by a single basis permutation. Let $\sigma:[d]\rightarrow[d]$ denote the mirror permutation with the action $\ell\mapsto d-\ell+1$ for each $\ell\in[d]$. Applying this permutation to the first $d$ elements of the basis of $\mathbb{R}^{2d}$, corresponding to the first summand, yields
$\mathtt{1}_{[\ell_{1}]}\oplus\mathtt{1}_{[\ell_{2}]}\ \mapsto\mathtt{1}_{\sigma([\ell_{1}])}\oplus\mathtt{1}_{[\ell_{2}]}=(0,\ldots 0,\underbrace{1,\ldots 1}_{\ell_{1}\text{times}}\mid\underbrace{1,\ldots 1,}_{\ell_{2}\text{times}}0,\ldots 0).$ (3.16)
The basic constraints are also given by vectors that have the consecutive ones property.
This is significant because a matrix with the consecutive ones property, also known as an interval matrix, is totally unimodular. That means the determinant of every one of its square submatrices equals $+1,0,$ or $-1$. This implies that if the alignment terms are integers, then the extremal points of $\mathcal{P}_{1}^{(k)}$ are integral, meaning their coordinates are integers (see Ch. 19 of Ref. [21]). From the basic constraints, this means that their coordinates are equal to 00 or $1$. Since a submatrix of a totally unimodular matrix is totally unimodular, the same is true for the extremal points of any polyhedron in $\mathcal{P}_{0}^{(k)}$ whose points satisfy a subset of the alignment constraints.
3.2 The meaning of alignment terms
We move on to discussing the meaning of alignment terms. To start, we observe that Fan’s maximum principle Eq. 2.2 implies the upper bound
$\alpha_{(\ell_{1},\ell_{2})}^{(k)}\leq\min(\ell_{1},k)+\min(\ell_{2},k),$ (3.17)
with equality if and only if $P_{\downarrow\ell_{1}}(A_{1})$ and $P_{\downarrow\ell_{2}}(A_{2})$ are perfectly aligned. We observe that Lidskii’s majorization relation (see Sec. III.4 of Ref. [4]) implies the lower bound
$\alpha_{(\ell_{1},\ell_{2})}^{(k)}\geq\min(k,\max(0,\ell_{1}+\ell_{2}-d))+\min(k,\ell_{1}+\ell_{2}).$ (3.18)
When $\ell_{1}+\ell_{2}\leq k$, the upper bound and lower bound coincide, implying the identity $\alpha_{(\ell_{1},\ell_{2})}^{(k)}=\ell_{1}+\ell_{2}$. Therefore, the cases where some information about the operators $A_{1}$ and $A_{2}$ is retained are the ones with $\ell_{1}+\ell_{2}>k$.
If the upper bound Eq. 3.17 is actually an equality, then the alignment constraint corresponding to the pair $(\ell_{1},\ell_{2})$ is redundant within $\mathcal{P}_{0}^{(k)}$. On the other hand, the slackness statement $\lceil\alpha_{(\ell_{1},\ell_{2})}^{(k)}\rceil<\min(\ell_{1},k)+\min(\ell_{2},k)$ implies an upper bound on $s_{k}(A_{1}+A_{2})$ with a combinatorial meaning. In words, $\lceil\alpha_{(\ell_{1},\ell_{2})}^{(k)}\rceil-k$ is the maximum number of eigenvalue pairs whose indices are in $[\ell_{1}]$ and $[\ell_{2}]$, respectively, that can appear (with coefficient $1$) together in the objective $(\lambda(A_{1})\oplus\lambda(A_{2}))^{T}x$. The remaining eigenvalue pairs have to be staggered. Notice, per the discussion in the previous paragraph, that the slackness statement implies $\ell_{1}+\ell_{2}>k$.
Proposition 3.3.
Let $m:=\lceil\alpha_{(\ell_{1},\ell_{2})}^{(k)}\rceil-k$ and suppose that it satisfies $m<\min(\ell_{1},k)+\min(\ell_{2},k)-k$. Denote the tuple of staggered eigenvalue sums
$\mu:=(\lambda_{i}(A_{1})+\lambda_{i-m+\ell_{2}}(A_{2}))_{i=m+1}^{k}\oplus(\lambda_{i-m+\ell_{1}}(A_{1})+\lambda_{i}(A_{2}))_{i=m+1}^{k}.$ (3.19)
Then $s_{k}(A_{1}+A_{2})\leq s_{m}(A_{1})+s_{m}(A_{2})+s_{k-m}(\mu)$.
Proof.
By Thm. 3.2, it suffices to show that the optimal value of the linear program
$\max\>(\lambda(A_{1})\oplus\lambda(A_{2}))^{T}x$ (3.20)
[…]
is $s_{m}(A_{1})+s_{m}(A_{2})+s_{k-m}(\mu)$. By total unimodularity, the extremal points of the feasible set of this program are integral (see Sec. 3.1). Hence, it suffices to maximize $(\lambda(A_{1})\oplus\lambda(A_{2}))^{T}(\mathtt{1}_{S_{1}}\oplus\mathtt{1}_{S_{2}})$ over set pairs $(S_{1},S_{2})\in[d]\times[d]$ that satisfy $|S_{1}|=|S_{2}|=k$ and $|S_{1}\cap[\ell_{1}]|+|S_{2}\cap[\ell_{2}]|\leq m+k.$
Given such a $k$-set pair $(S_{1},S_{2})$, we may assume without loss of generality that $S_{1}\cap[\ell_{1}]=[r_{1}]$ and $S_{2}\cap[\ell_{2}]=[r_{2}]$ where $r_{1}:=|S_{1}\cap[\ell_{1}]|$ and $r_{2}:=|S_{2}\cap[\ell_{2}]|$. Otherwise, the intersections $S_{1}\cap[\ell_{1}]$ and $S_{2}\cap[\ell_{2}]$ may be adjusted to improve the objective value. By the same token, we may assume that $S_{1}\cap[\ell_{1}]^{c}=\{\ell_{1}+1,\ldots,\ell_{1}+k-r_{1}\}$ and $S_{2}\cap[\ell_{2}]^{c}=\{\ell_{2}+1,\ldots,\ell_{2}+k-r_{2}\}$. Lastly, we argue that the alignment constraint may be assumed to be satisfied with equality. If $|S_{1}\cap[\ell_{1}]|+|S_{2}\cap[\ell_{2}]|\neq m+k$, then $S_{1}$ may be adjusted to improve the objective value by replacing elements in $S_{1}\cap[\ell_{1}]^{c}$ with elements in $[\ell_{1}]$ until the constraint is satisfied with equality or $S_{1}\cap[\ell_{1}]^{c}$ becomes empty. If the latter happens before the former, $S_{2}$ may be adjusted to improve the objective value by replacing elements in $S_{2}\cap[\ell_{2}]^{c}$ with elements in $[\ell_{2}]$ until the constraint is satisfied with equality. Hence, without loss of generality, we assume that $r_{1}+r_{2}=k+m$. From the fact that $\max(r_{1},r_{2})\leq k$, we conclude that $m\leq\min(r_{1},r_{2})$.
With these assumptions on $(S_{1},S_{2})$, the objective function is
$(\lambda(A_{1})\oplus\lambda(A_{2}))^{T}(\mathtt{1}_{S_{1}}\oplus\mathtt{1}_{S_{2}})$ $=s_{m}(A_{1})+s_{m}(A_{2})$ (3.21)
$+\sum_{i=m+1}^{r_{1}}\lambda_{i}(A_{1})+\sum_{i=1}^{r_{1}-m}\lambda_{i+\ell_{2}}(A_{2})$
$+\sum_{i=1}^{k-r_{1}}\lambda_{i+\ell_{1}}(A_{1})+\sum_{i=m+1}^{k+m-r_{1}}\lambda_{i}(A_{2}).$
We rewrite the sums in the second and third lines.
$\sum_{i=m+1}^{r_{1}}\lambda_{i}(A_{1})+\sum_{i=1}^{r_{1}-m}\lambda_{i+\ell_{2}}(A_{2})=\sum_{i=m+1}^{r_{1}}\lambda_{i}(A_{1})+\lambda_{i-m+\ell_{2}}(A_{2}),$ (3.22)
$\sum_{i=1}^{k-r_{1}}\lambda_{i+\ell_{1}}(A_{1})+\sum_{i=m+1}^{m+k-r_{1}}\lambda_{i}(A_{2})=\sum_{i=m+1}^{m+k-r_{1}}\lambda_{i-m+\ell_{1}}(A_{1})+\lambda_{i}(A_{2}).$
Then, by the definition of $\mu$, it holds that
$\sum_{i=m+1}^{r_{1}}(\lambda_{i}(A_{1})+\lambda_{i-m+\ell_{2}}(A_{2}))$ $+\sum_{i=m+1}^{m+k-r_{1}}(\lambda_{i-m+\ell_{1}}(A_{1})+\lambda_{i}(A_{2}))$ (3.23)
$\leq s_{k-m}(\mu).$
Equality is realized by maximizing over $r_{1}$. ∎
3.3 Simultaneously diagonalizable summands
In this subsection, we outline a strategy for proving that the linear programming upper bound given by Thm. 3.2 is tight in the case of simultaneously diagonalizable summands. We complete the proof in App. A.
Let $D_{1}$ and $D_{2}$ be two simultaneously diagonalizable self-adjoint operators on $\mathbb{C}^{d}$. Without loss of generality, we assume that $D_{1}$ and $D_{2}$ are diagonal in the standard basis $\{q_{i}\}_{i=1}^{d}$. Furthermore, we take $V_{\bullet}(D_{1})$ and $V_{\bullet}(D_{2})$ to be complete flags corresponding to a shared diagonal spectral decomposition of $D_{1}$ and $D_{2}$. For each $i\in[d]$, we denote $\tilde{\lambda}_{i}(D_{1}):=q_{i}^{*}D_{1}q_{i}$ and $\tilde{\lambda}_{i}(D_{2}):=q_{i}^{*}D_{2}q_{i}$. To capture the order of the eigenvalues, we define the sets
$\Omega^{1}_{\ell}:=\{i\in[d]\mid P_{\downarrow\ell}(D_{1})q_{i}=q_{i}\},\Omega^{2}_{\ell}:=\{i\in[d]\mid P_{\downarrow\ell}(D_{2})q_{i}=q_{i}\}.$ (3.24)
for each $\ell\in[d]$. These are the sets of coordinates of the largest $\ell$ values in $\tilde{\lambda}(D_{1}):=(\tilde{\lambda}_{i}(D_{1}))_{i=1}^{d}$ and $\tilde{\lambda}(D_{2}):=(\tilde{\lambda}_{i}(D_{2}))_{i=1}^{d}$, respectively. We use the two subset chains $\{\Omega^{1}_{\ell_{1}}\}_{\ell_{1}=1}^{d}$ and $\{\Omega^{2}_{\ell_{2}}\}_{\ell_{2}=1}^{d}$ to define two total orders, $\geq^{1}$ and $\geq^{2}$, on $[d]$. For $x,x^{\prime}\in[d]$, we write $x\geq^{1}x^{\prime}$ if
$\min\{\ell\mid x\in\Omega^{1}_{\ell}\}\geq\min\{\ell\mid x^{\prime}\in\Omega^{1}_{\ell}\}.$ (3.25)
Similarly, we write $x\geq^{2}x^{\prime}$ if
$\min\{\ell\mid x\in\Omega^{2}_{\ell}\}\geq\min\{\ell\mid x^{\prime}\in\Omega^{2}_{\ell}\}.$ (3.26)
Observe that $x\geq^{1}x^{\prime}$ implies $\tilde{\lambda}_{x}(D_{1})\leq\tilde{\lambda}_{x^{\prime}}(D_{1})$ and $x\geq^{2}x^{\prime}$ implies $\tilde{\lambda}_{x}(D_{2})\leq\tilde{\lambda}_{x^{\prime}}(D_{2})$.
Now, we compute the alignment terms of $D_{1}$ and $D_{2}$.
Lemma 3.4.
For each $(\ell_{1},\ell_{2})\in[d]\times[d]$, it holds that
$\alpha_{(\ell_{1},\ell_{2})}^{(k)}(V_{\bullet}(D_{1}),V_{\bullet}(D_{2}))=\min(k,|\Omega^{1}_{\ell_{1}}\cap\Omega^{2}_{\ell_{2}}|)+\min(k,|\Omega^{1}_{\ell_{1}}\cup\Omega^{2}_{\ell_{2}}|).$ (3.27)
Proof.
Recall that
[…]
$=s_{k}(P_{\downarrow\ell_{1}}(D_{1})+P_{\downarrow\ell_{2}}(D_{2}))$ (3.29)
$=s_{k}(\sum_{i\in\Omega^{1}_{\ell_{1}}}q_{i}q_{i}^{*}+\sum_{i\in\Omega^{2}_{\ell_{2}}}q_{i}q_{i}^{*}).$ (3.30)
If $k\leq|\Omega^{1}_{\ell_{1}}\cap\Omega^{2}_{\ell_{2}}|$, then it equals $2k$. If $|\Omega^{1}_{\ell_{1}}\cap\Omega^{2}_{\ell_{2}}|<k<|\Omega^{1}_{\ell_{1}}\cup\Omega^{2}_{\ell_{2}}|$, then it equals $2|\Omega^{1}_{\ell_{1}}\cap\Omega^{2}_{\ell_{2}}|+k-|\Omega^{1}_{\ell_{1}}\cap\Omega^{2}_{\ell_{2}}|=|\Omega^{1}_{\ell_{1}}\cap\Omega^{2}_{\ell_{2}}|+k$. Lastly, if $k\geq|\Omega^{1}_{\ell_{1}}\cup\Omega^{2}_{\ell_{2}}|$, then it equals $|\Omega^{1}_{\ell_{1}}|+|\Omega^{2}_{\ell_{2}}|$ which equals $|\Omega^{1}_{\ell_{1}}\cap\Omega^{2}_{\ell_{2}}|+|\Omega^{1}_{\ell_{1}}\cup\Omega^{2}_{\ell_{2}}|$. ∎
Since the alignment terms are integers, the coordinates of the extremal points of $\mathcal{P}_{1}^{(k)}(V_{\bullet}(D_{1}),V_{\bullet}(D_{2}))$ are in $\{0,1\}$. Therefore, the linear programming upper bound $u_{k}(D_{1},D_{2})$ equals the optimal value of the following combinatorial optimization problem:
$\max\;(\tilde{\lambda}(D_{1})\oplus\tilde{\lambda}(D_{2}))^{T}(\mathtt{1}_{S_{1}}\oplus\mathtt{1}_{S_{2}})$ (3.31)
over subsets pairs $(S_{1},S_{2})\subseteq[d]\times[d]$ satisfying $|S_{1}|=|S_{2}|=k$ and for all $(\ell_{1},\ell_{2})\in[d]\times[d]$,
$|S_{1}\cap\Omega^{1}_{\ell_{1}}|+|S_{2}\cap\Omega^{2}_{\ell_{2}}|\leq\min(k,|\Omega^{1}_{\ell_{1}}\cap\Omega^{2}_{\ell_{2}}|)+\min(k,|\Omega^{1}_{\ell_{1}}\cup\Omega^{2}_{\ell_{2}}|).$ (3.32)
We call a subset pair $(S_{1},S_{2})$ that satisfies all of these constraints a feasible pair. Notice that the objective here differs from the one in Thm. 3.2 by a relabeling of the coordinates of the linear form.
Next, we observe that $s_{k}(D_{1}+D_{2})$ equals
$\max_{S\subseteq[d],|S|=k}\sum_{i\in S}q_{i}^{*}(D_{1}+D_{2})q_{i}=\max_{S\subseteq[d],|S|=k}\sum_{i\in S}\tilde{\lambda}_{i}(D_{1})+\tilde{\lambda}_{i}(D_{2})$ (3.33)
which can be rewritten as follows
$\max_{(S,S)\subseteq[d]\times[d],|S|=k}\;(\tilde{\lambda}(D_{1})\oplus\tilde{\lambda}(D_{2}))^{T}(\mathtt{1}_{S}\oplus\mathtt{1}_{S}).$ (3.34)
We call set pairs of the form $(S,S)$ symmetric. Therefore, to prove that $u_{k}(D_{1},D_{2})=s_{k}(D_{1}+D_{2})$, it suffices to show that for an arbitrary feasible pair $(S_{1},S_{2})$, there exists a feasible symmetric pair $(S,S)$ such that the objective Eq. 3.31 at $(S,S)$ is greater than or equal to the objective at $(S_{1},S_{2})$.
Theorem 3.5.
$u_{k}(D_{1},D_{2})=s_{k}(D_{1}+D_{2})$.
Proof.
Deferred to appendix App. A. ∎
4 Applications to sums of tensor products
We use theorems Thm. 3.2 and Thm. 3.5 to prove our main theorem Thm. 1.3.
Fix two positive integers $d_{B}$ and $d_{C}$. Let $B_{1},B_{2}$ be positive semi-definite operators on $\mathbb{C}^{d_{B}}$ and $C_{1},C_{2}$ be positive semi-definite operators on $\mathbb{C}^{d_{C}}$. The goal is to prove the majorization relation
$B_{1}\otimes C_{1}+B_{2}\otimes C_{2}\preceq B_{1}^{\downarrow}\otimes C_{1}^{\downarrow}+B_{2}^{\downarrow}\otimes C_{2}^{\downarrow}.$ (4.1)
We start with an example showing that the assumption of positive semi-definiteness cannot be relaxed in general.
Example 4.1.
Consider the following operators on $\mathbb{C}^{2}$: $\tilde{B}_{1}:=q_{1}q_{1}^{*}+q_{2}q_{2}^{*}=:-\tilde{B}_{2},\tilde{C}_{1}:=q_{1}q_{1}^{*},$ and $\tilde{C}_{2}:=q_{2}q_{2}^{*}$. The spectrum of the sum $\tilde{B}_{1}\otimes\tilde{C}_{1}+\tilde{B}_{2}\otimes\tilde{C}_{2}$ is $(1,1,-1,-1)$. On the other hand, the aligned sum $\tilde{B}_{1}^{\downarrow}\otimes\tilde{C}_{1}^{\downarrow}+\tilde{B}_{2}^{\downarrow}\otimes\tilde{C}_{2}^{\downarrow}$ is null, hence its spectrum is $(0,0,0,0)$.
4.1 Ordering the spectrum
Let $\tilde{B}$ be a positive semi-definite operator on $\mathbb{C}^{d_{B}}$ and $\tilde{C}$ be a positive semi-definite operator on $\mathbb{C}^{d_{C}}$. Their tensor product $\tilde{B}\otimes\tilde{C}$ has the spectral decomposition
$\sum_{(i,j)\in[d_{B}]\times[d_{C}]}\lambda_{i}(\tilde{B})\lambda_{j}(\tilde{C})\Pi_{i}(\tilde{B})\otimes\Pi_{j}(\tilde{C}).$ (4.2)
The order of the eigenvalues of this operator is not determined entirely by the orders of the eigenvalues of $\tilde{B}$ and $\tilde{C}$. In general, for $(i,j),(i^{\prime},j^{\prime})\in[d_{B}]\times[d_{C}]$, whether the inequality
$\lambda_{i}(\tilde{B})\lambda_{j}(\tilde{C})\leq\lambda_{i^{\prime}}(\tilde{B})\lambda_{j^{\prime}}(\tilde{C})$ (4.3)
holds or not depends on the eigenvalues in question, not only their order. Here is what we can conclude however. Since the eigenvalues are non-negative, if $i\geq i^{\prime}$ and $j\geq j^{\prime}$, then the inequality holds. That is to say, the orders of the eigenvalues of $\tilde{B}$ and $\tilde{C}$ induce a partial order on the eigenvalues of $\tilde{B}\otimes\tilde{C}$, namely the product order. We denote this relation by $\geq^{\times}$.
To account for the remaining order relations implied by the spectra of $\tilde{B}$ and $\tilde{C}$, we associate with $\tilde{B}\otimes\tilde{C}$ a chain of downward closed sets in $([d_{B}]\times[d_{C}],\geq^{\times})$
$\Upsilon_{1}=\{(1,1)\}\subset\Upsilon_{2}\subset\cdots\subset\Upsilon_{d_{B}d_{C}}=[d_{B}]\times[d_{C}]$ (4.4)
such that for each $\ell\in[d_{B}d_{C}]$, we have
$V_{\downarrow\ell}(\tilde{B}\otimes\tilde{C})=\langle\{\xi_{i}(\tilde{B})\otimes\xi_{j}(\tilde{C})\mid(i,j)\in\Upsilon_{\ell}\}\rangle.$ (4.5)
Observe that this implies $|\Upsilon_{\ell}|=\ell$.
That such a chain exists can be proven by induction. The chain in Eq. 4.4 induces a total order on $[d_{B}]\times[d_{C}]$. For $(i,j),(i^{\prime},j^{\prime})\in[d_{B}]\times[d_{C}]$, $(i,j)$ is greater than or equal to $(i^{\prime},j^{\prime})$ in this order if
$\min\{\ell\mid(i,j)\in\Upsilon_{\ell}\}\geq\min\{\ell\mid(i^{\prime},j^{\prime})\in\Upsilon_{\ell}\}.$ (4.6)
Observe that if this holds, then $\lambda_{i}(\tilde{B})\lambda_{j}(\tilde{C})\leq\lambda_{i^{\prime}}(\tilde{B})\lambda_{j^{\prime}}(\tilde{C})$. Fig. 1 depicts a Hasse diagram for $([2]\times[2],\geq^{\times})$ along with a total order induced by a chain of downward closed sets. This is the order of the eigenvalues of $\tilde{B}\otimes\tilde{C}$ when, for example, $\tilde{B}=3q_{1}q_{1}^{*}+q_{2}q_{2}^{*}$ and $\tilde{C}=3q_{1}q_{1}^{*}+2q_{2}q_{2}^{*}$.
Let $\{\Upsilon^{1}_{\ell}\}_{\ell=1}^{d_{B}d_{C}}$ and $\{\Upsilon^{2}_{\ell}\}_{\ell=1}^{d_{B}d_{C}}$ denote the two chains of downward closed sets associated with $B_{1}\otimes C_{1}$ and $B_{2}\otimes C_{2}$, respectively.
$(2,2)$$(1,2)$$(2,1)$$(1,1)$
Figure 1: A Hasse diagram for $([2]\times[2],\geq^{\times})$. The curved arrow indicates a total order on $[2]\times[2]$ induced by the chain $\{(1,1)\}\subset\{(1,1),(1,2)\}\subset\{(1,1),(1,2),(2,1)\}\subset\{(1,1),(1,2),(2,1),(2,2)\}$. According to this total order, $(2,1)$ is greater than or equal to $(1,2)$, though the two are incomparable in the product order.
4.2 Bounding the alignment terms
To bound the alignment term $\alpha_{(\ell_{1},\ell_{2})}^{(k)}(V_{\bullet}(B_{1}\otimes C_{1}),V_{\bullet}(B_{2}\otimes C_{2}))$, we examine spaces spanned by product vectors indexed by elements of downward closed sets. Upon our examination, we find a generalization of a well known fact about subspace intersections. Specifically, that for any two orthonormal bases $\{f_{i}\}_{i=1}^{d}$ and $\{g_{i}\}_{i=1}^{d}$ for $\mathbb{C}^{d}$, and two downward closed sets $Y,Y^{\prime}\subseteq[d]$, it holds that
$\dim(\langle\{f_{i}\}_{i\in Y}\rangle\cap\langle\{g_{i}\}_{i\in Y^{\prime}}\rangle^{\perp})\geq|Y\setminus Y^{\prime}|.$ (4.7)
Let $\{f_{i}^{(B)}\}_{i=1}^{d_{B}}$ and $\{g_{i}^{(B)}\}_{i=1}^{d_{B}}$ be two orthonormal bases for $\mathbb{C}^{d_{B}}$, and $\{f_{j}^{(C)}\}_{j=1}^{d_{C}}$ and $\{g_{j}^{(C)}\}_{j=1}^{d_{C}}$ be two orthonormal bases for $\mathbb{C}^{d_{C}}$. For the sake of readability, we denote for $i\in[d_{B}]$,
$F^{(B)}_{\downarrow i}:=\langle\{f^{(B)}_{1},\ldots,f^{(B)}_{i}\}\rangle\,,\,G^{(B)}_{\uparrow i}:=\langle\{g_{i}^{(B)},\ldots,g_{d_{B}}^{(B)}\}\rangle,$ (4.8)
for $j\in[d_{C}]$,
$F^{(C)}_{\downarrow j}:=\langle\{f^{(C)}_{1},\ldots,f^{(C)}_{j}\}\rangle\,,\,G^{(C)}_{\uparrow j}:=\langle\{g^{(C)}_{j},\ldots,g^{(C)}_{d_{C}}\}\rangle,$ (4.9)
and, finally, for a subset $S\subseteq[d_{B}]\times[d_{C}]$,
$F_{S}:=\langle\{f_{i}^{(B)}\otimes f_{j}^{(C)}\mid(i,j)\in S\}\rangle,$ (4.10)
$G_{S}:=\langle\{g_{i}^{(B)}\otimes g_{j}^{(C)}\mid(i,j)\in S\}\rangle.$ (4.11)
Proposition 4.2.
Let $\Upsilon$ and $\Upsilon^{\prime}$ be two downward closed sets in $([d_{B}]\times[d_{C}],\geq^{\times})$. Then
$\dim(F_{\Upsilon}\cap{G_{\Upsilon^{\prime}}}^{\perp})\geq|\Upsilon\setminus\Upsilon^{\prime}|.$ (4.12)
Proof.
Observe that ${G_{\Upsilon^{\prime}}}^{\perp}=G_{{\Upsilon^{\prime}}^{c}}$. Denote the difference $\Delta:=\Upsilon\setminus\Upsilon^{\prime}$. Since $\downarrow\Delta\subseteq\Upsilon$ and $\uparrow\Delta\subseteq{\Upsilon^{\prime}}^{c}$, it suffices to prove $\dim(F_{\downarrow\Delta}\cap G_{\uparrow\Delta})\geq|\Delta|.$ We do this by showing the existence of a subspace whose dimension is not so large, yet the dimensions of its intersections with $F_{\downarrow\Delta}$ and $G_{\uparrow\Delta}$ are relatively large.
As the difference of two downward closed sets, $\Delta$ is interval-like. That is, if $x,y\in\Delta$ and $x\leq^{\times}z\leq^{\times}y$, then $z\in\Delta$. This implies ${\downarrow\Delta}\cap{\uparrow\Delta}=\Delta$. Let $\Delta_{1}:=\{i_{1},\ldots,i_{T}\}\subseteq[d_{B}],\Delta_{2}:=\{j_{1},\ldots,j_{L}\}\subseteq[d_{C}]$ denote the sets of first coordinates and second coordinates, respectively, of the elements of $\Delta$. Let $(\tilde{i},\tilde{j})$ be a point in $\Delta_{1}\times\Delta_{2}$. There exists $j^{\prime}\in[d_{C}]$ such that $(\tilde{i},j^{\prime})\in\Delta$. If $j^{\prime}\geq\tilde{j}$, then $(\tilde{i},\tilde{j})\in{\downarrow\Delta}$. Otherwise, $(\tilde{i},\tilde{j})\in{\uparrow\Delta}$. Therefore, the inclusion $\Delta_{1}\times\Delta_{2}\subseteq{\downarrow\Delta}\cup{\uparrow\Delta}$ holds. See Fig. 2 for an illustrative example.
$(2,2)$$(1,2)$$(2,1)$$(1,1)$$\mathbf{(1,3)}$$\mathbf{(3,1)}$$(2,3)$$(3,2)$$(3,3)$
Figure 2: A Hasse diagram for $([3]\times[3],\geq^{\times})$. In this example, the set $\Delta$ is $\{(1,3),(3,1)\}$. The downward closure ${\downarrow\Delta}=\{(1,3),(1,2),(1,1),(2,1),(3,1)\}$ is indicated by the downward facing arrows, while the upward closure ${\uparrow\Delta}=\{(1,3),(2,3),(3,3),(3,2),(3,1)\}$ is indicated by the upward facing arrows. In this example, the product $\Delta_{1}\times\Delta_{2}$ is $\{(1,3),(1,1),(3,1),(3,3)\}$.
By a dimension counting argument due to Amir-Moéz (see Lem 6.1 in Ref. [2] and Sec. III.3 in Ref. [4]), there exist an orthonormal set $\{\tilde{f}^{(B)}_{i_{t}}\}_{t=1}^{T}$ satisfying $\tilde{f}^{(B)}_{i_{t}}\in U^{(B)}_{\downarrow i_{t}}$ for each $t\in[T]$ and an orthonormal set $\{\tilde{g}^{(B)}_{i_{t}}\}_{t=1}^{T}$ satisfying $\tilde{g}^{(B)}_{i_{t}}\in G^{(B)}_{\uparrow i_{t}}$ for each $t\in[T]$ such that
$\langle\{\tilde{f}^{(B)}_{i_{t}}\}_{t=1}^{T}\rangle=\langle\{\tilde{g}^{(B)}_{i_{t}}\}_{t=1}^{T}\rangle=:W^{(1)}$ (4.13)
By the same token, there exist an orthonormal set $\{\tilde{f}^{(C)}_{j_{\ell}}\}_{\ell=1}^{L}$ satisfying $\tilde{f}^{(C)}_{j_{\ell}}\in F^{(C)}_{\downarrow j_{\ell}}$ for each $\ell\in[L]$ and an orthonormal set $\{\tilde{g}^{(C)}_{j_{\ell}}\}_{\ell=1}^{L}$ satisfying $\tilde{g}^{(C)}_{j_{\ell}}\in G^{(C)}_{\uparrow j_{\ell}}$ for each $\ell\in[L]$ such that
$\langle\{\tilde{f}^{(C)}_{j_{\ell}}\}_{\ell=1}^{L}\rangle=\langle\{\tilde{g}^{(C)}_{j_{\ell}}\}_{\ell=1}^{L}\rangle=:W^{(2)}.$ (4.14)
The subspace we seek is the tensor product $W:=W^{(1)}\otimes W^{(2)}$.
For $(i,j)\in{\downarrow\Delta}$, the inclusion $\downarrow\{(i,j)\}\subseteq{\downarrow\Delta}$ implies $F_{\downarrow i}^{(B)}\otimes F_{\downarrow j}^{(C)}=F_{\downarrow\{(i,j)\}}\leq F_{\downarrow\Delta}$. Hence, for any $t\in[T],\ell\in[L]$ such that $(i_{t},j_{\ell})\in{\downarrow\Delta}$, the membership $\tilde{f}^{(B)}_{i_{t}}\otimes\tilde{f}^{(C)}_{j_{\ell}}\in F_{\downarrow\Delta}$ holds. This implies
$\dim(F_{\downarrow\Delta}\cap W)\geq|(\Delta_{1}\times\Delta_{2})\cap{\downarrow\Delta}|.$ (4.15)
Similar considerations for the other basis yield
$\dim(G_{\uparrow\Delta}\cap W)\geq|(\Delta_{1}\times\Delta_{2})\cap{\uparrow\Delta}|.$ (4.16)
Putting it altogether, we obtain the following estimate.
$\dim(F_{\downarrow\Delta}\cap G_{\uparrow\Delta})$ $\geq\dim((F_{\downarrow\Delta}\cap W)\cap(G_{\uparrow\Delta}\cap W))$
$\geq\dim(F_{\downarrow\Delta}\cap W)+\dim(G_{\uparrow\Delta}\cap W)-\dim(W)$
$\geq|(\Delta_{1}\times\Delta_{2})\cap{\downarrow\Delta}|+|(\Delta_{1}\times\Delta_{2})\cap{\uparrow\Delta}|-|\Delta_{1}\times\Delta_{2}|$
$=|\Delta|,$ (4.17)
where for the equality we used inclusion-exclusion and the facts $\Delta_{1}\times\Delta_{2}\subseteq{\downarrow\Delta}\cup{\uparrow\Delta}$ and ${\downarrow\Delta}\cap{\uparrow\Delta}=\Delta\subseteq\Delta_{1}\times\Delta_{2}$. ∎
With this in hand, we can obtain tight upper bounds for the alignment terms of tensor product operators.
Proposition 4.3.
For each pair $(\ell_{1},\ell_{2})\in[d_{B}d_{C}]\times[d_{B}d_{C}]$,
$\alpha_{(\ell_{1},\ell_{2})}^{(k)}(V_{\bullet}(B_{1}\otimes C_{1}),V_{\bullet}(B_{2}\otimes C_{2}))\leq\alpha_{(\ell_{1},\ell_{2})}^{(k)}(V_{\bullet}(B_{1}^{\downarrow}\otimes C_{1}^{\downarrow}),V_{\bullet}(B_{2}^{\downarrow}\otimes C_{2}^{\downarrow})).$
Proof.
By the previous proposition, $V_{\downarrow\ell_{1}}(B_{1}\otimes C_{1})$ contains a subspace of dimension $|\Upsilon^{1}_{\ell_{1}}\setminus\Upsilon^{2}_{\ell_{2}}|$ that is orthogonal to $V_{\downarrow\ell_{2}}(B_{2}\otimes C_{2})$. By the same token, $V_{\downarrow\ell_{2}}(B_{2}\otimes C_{2})$ contains a subspace of dimension $|\Upsilon^{2}_{\ell_{2}}\setminus\Upsilon^{1}_{\ell_{1}}|$ that is orthogonal to $V_{\downarrow\ell_{1}}(B_{1}\otimes C_{1})$. Hence, the multiplicity of $1$ in the spectrum of $P_{\downarrow\ell_{1}}(B_{1}\otimes C_{1})+P_{\downarrow\ell_{2}}(B_{2}\otimes C_{2})$ is at least $|\Upsilon^{1}_{\ell_{1}}\setminus\Upsilon^{2}_{\ell_{2}}|+|\Upsilon^{2}_{\ell_{2}}\setminus\Upsilon^{1}_{\ell_{1}}|$. The sum of remainder of the eigenvalues is $2|\Upsilon^{(B)}_{\ell_{1}}\cap\Upsilon^{(C)}_{\ell_{2}}|$, and so
$\lambda(P_{\downarrow\ell_{1}}(B_{1}\otimes C_{1})+P_{\downarrow\ell_{2}}(B_{2}\otimes C_{2}))\preceq(2,\ldots,2,1,\ldots,1,0,\ldots,0),$ (4.18)
where on the right hand side the multiplicity of $2$ is $|\Upsilon^{(B)}_{\ell_{1}}\cap\Upsilon^{(C)}_{\ell_{2}}|$, and the multiplicity of $1$ is $|\Upsilon^{(B)}_{\ell_{1}}\setminus\Upsilon^{(C)}_{\ell_{2}}|+|\Upsilon^{(C)}_{\ell_{2}}\setminus\Upsilon^{(B)}_{\ell_{1}}|$. The proposition follows by observing that the right hand side of Eq. 4.18 equals $\lambda(P_{\downarrow\ell_{1}}(B_{1}^{\downarrow}\otimes C_{1}^{\downarrow})+P_{\downarrow\ell_{2}}(B_{2}^{\downarrow}\otimes C_{2}^{\downarrow}))$.∎
We are now ready to prove Thm. 1.3.
Proof of Thm. 1.3.
An immediate corollary of Prop. 4.3 is that
$\mathcal{P}_{1}^{(k)}(V_{\bullet}(B_{1}\otimes C_{1}),V_{\bullet}(B_{2}\otimes C_{2}))\subseteq\mathcal{P}_{1}^{(k)}(V_{\bullet}(B_{1}^{\downarrow}\otimes C_{1}^{\downarrow}),V_{\bullet}(B_{2}^{\downarrow}\otimes C_{2}^{\downarrow})).$ (4.19)
This implies that the corresponding linear programming bounds satisfy
$u_{k}(B_{1}\otimes C_{1},B_{2}\otimes C_{2})\leq u_{k}(B_{1}^{\downarrow}\otimes C_{1}^{\downarrow},B_{2}^{\downarrow}\otimes C_{2}^{\downarrow}).$ (4.20)
Since $B_{1}^{\downarrow}\otimes C_{1}^{\downarrow}$ and $B_{2}^{\downarrow}\otimes C_{2}^{\downarrow}$ are simultaneously diagonalizable, it follows from Thm. 3.5 that
$s_{k}(B_{1}^{\downarrow}\otimes C_{1}^{\downarrow}+B_{2}^{\downarrow}\otimes C_{2}^{\downarrow})=u_{k}(B_{1}^{\downarrow}\otimes C_{1}^{\downarrow},B_{2}^{\downarrow}\otimes C_{2}^{\downarrow}).$ (4.21)
The theorem follows from the fact that
$s_{k}(B_{1}\otimes C_{1}+B_{2}\otimes C_{2})\leq u_{k}(B_{1}\otimes C_{1},B_{2}\otimes C_{2}).$ (4.22)
4.3 Resolving the $2$-letter spin alignment conjecture
We use Eq. 1.9 to resolve the $n=2$ case of the spin alignment conjecture.
Proposition 4.4.
Let $M$ denote a quantum state on $\mathbb{C}^{d}$ and $p$ denote a probability distribution on the subsets of $\{1,2\}$. Then for every pure quantum state tuple $(1,\psi_{\{1\}},\psi_{\{2\}},\psi_{\{1,2\}})$ and every unit vector $v_{M}$ satisfying $Mv_{M}=\lambda_{1}(M)v_{M}$, it holds that
$\sum_{J\subseteq\{1,2\}}p_{J}\psi_{J}\otimes M^{\otimes J^{c}}\preceq\sum_{J\subseteq\{1,2\}}p_{J}({v_{M}{v_{M}}^{*}})^{\otimes J}\otimes M^{\otimes J^{c}}.$ (4.23)
Proof.
Since the two operators $p_{\varnothing}M\otimes M+p_{\{1\}}v_{M}v_{M}^{*}\otimes M+p_{\{2\}}M\otimes v_{M}v_{M}^{*}$ and $p_{\{1,2\}}v_{M}v_{M}^{*}\otimes v_{M}v_{M}^{*}$ are perfectly aligned, it suffices to prove that $p_{\varnothing}M\otimes M+p_{\{1\}}v_{M}v_{M}^{*}\otimes M+p_{\{2\}}M\otimes v_{M}v_{M}^{*}$ majorizes $p_{\varnothing}M\otimes M+p_{\{1\}}\psi_{1}\otimes M+p_{\{2\}}M\otimes\psi_{2}$.
Combining the first two terms of the latter and using Eq. 1.9 yields
$(p_{\varnothing}M+p_{\{1\}}\psi_{1})\otimes M$ $+p_{\{2\}}M\otimes\psi_{2}\preceq$ (4.24)
$(p_{\varnothing}M+p_{\{1\}}\psi_{1})^{\downarrow}\otimes M^{\downarrow}+p_{\{2\}}M^{\downarrow}\otimes\psi_{2}^{\downarrow}.$
The right hand side of this relation is unitarily equivalent to
$(p_{\varnothing}M+p_{\{1\}}\psi_{1})^{\downarrow}\otimes M+p_{\{2\}}M^{\downarrow}\otimes v_{M}v_{M}^{*}.$ (4.25)
Define $\tilde{M}:=M-\lambda_{1}(M)v_{M}v_{M}^{*}$. Expanding the factor $M$ in the first term of the operator in Eq. 4.25 yields
$(p_{\varnothing}M+p_{\{1\}}\psi_{1})^{\downarrow}\otimes\tilde{M}+(\lambda_{1}(M)(p_{\varnothing}M+p_{\{1\}}\psi_{1})^{\downarrow}+p_{\{2\}}M^{\downarrow})\otimes v_{M}v_{M}^{*}.$ (4.26)
Since the two terms are orthogonal, we can treat them separately (compare with Prop III.8 in Ref. [1]). Fan’s majorization relation Eq. 1.8 and, for example, Theorem 2.2 in Ref. [6] imply
$(p_{\varnothing}M+p_{\{1\}}\psi_{1})^{\downarrow}\otimes\tilde{M}\preceq(p_{\varnothing}M^{\downarrow}+p_{\{1\}}\psi_{1}^{\downarrow})\otimes\tilde{M}.$ (4.27)
The right hand side of this relation is unitarily equivalent to $(p_{\varnothing}M+p_{\{1\}}v_{M}v_{M}^{*})\otimes\tilde{M}$. By the same token, it holds that
$(\lambda_{1}(M)(p_{\varnothing}M$ $+p_{\{1\}}\psi_{1})^{\downarrow}+p_{\{2\}}M^{\downarrow})\otimes v_{M}v_{M}^{*}$ (4.28)
$\preceq(\lambda_{1}(M)(p_{\varnothing}M^{\downarrow}+p_{\{1\}}\psi_{1}^{\downarrow})+p_{\{2\}}M^{\downarrow})\otimes v_{M}v_{M}^{*}.$
The right hand side of this relation is unitarily equivalent to
$(\lambda_{1}(M)(p_{\varnothing}M+p_{\{1\}}v_{M}v_{M}^{*})+p_{\{2\}}M)\otimes v_{M}v_{M}^{*}.$ (4.29)
Therefore, by the reduction in Sec. V of Ref. [14], the coherent information of any platypus channel is additive to the $2$-letter level.
Corollary 4.5.
If $\mathcal{M}$ is a platypus channel, then $I^{\text{(coh)}}({\mathcal{M}}^{\otimes 2})=2I^{\text{(coh)}}({\mathcal{M}})$.
5 Concluding remarks
The obstacle to generalizing Eq. 1.9 to three or more factors is that the proof of Prop. 4.2 does not apply to cases with more than two factors.
The set of triples
$\{(\lambda(A_{1}),\lambda(A_{2}),\lambda(A_{1}+A_{2}))\mid A_{1},A_{2}\;\text{self-adjoint on}\;\mathbb{C}^{d_{B}}\otimes\mathbb{C}^{d_{C}}\}$ (5.1)
is a polyhedron that is completely described by Horn’s inequalities [5]. These include the inequalities given by Fan’s majorization relation. The new majorization relation Eq. 1.9 gives new inequalities that account for the tensor product structure of the summands. Majorization relations involving tensor products feature in the theory of catalytic majorization [3].
Finally, here is an approach to proving Eq. 1.9 that does not work. Observe that the majorization relation
$B_{1}^{\downarrow}\otimes C_{1}+B_{2}^{\downarrow}\otimes C_{2}\preceq B_{1}^{\downarrow}\otimes C_{1}^{\downarrow}+B_{2}^{\downarrow}\otimes C_{2}^{\downarrow}$ (5.2)
holds. Since majorization is transitive, Eq. 1.9 would be implied by the majorization relation
$B_{1}\otimes C_{1}+B_{2}\otimes C_{2}\preceq^{?}B_{1}^{\downarrow}\otimes C_{1}+B_{2}^{\downarrow}\otimes C_{2}.$ (5.3)
However, this latter majorization relation does not hold in general.
Example 5.1.
Let $d_{B}=d_{C}=2$, and denote $e:=\frac{1}{\sqrt{2}}(q_{1}+q_{2})$. Let $\tilde{B}_{1}:=\tilde{C}_{1}:=ee^{*}$, $\tilde{B}_{2}=2q_{1}q_{1}^{*}+q_{2}q_{2}^{*}$ and $\tilde{C}_{2}=q_{1}q_{1}^{*}$. Then
$\tilde{B}_{1}\otimes\tilde{C_{1}}+\tilde{B}_{2}\otimes\tilde{C}_{2}=ee^{*}\otimes ee^{*}+2q_{1}q_{1}^{*}\otimes q_{1}q_{1}^{*}+q_{2}q_{2}^{*}\otimes q_{1}q_{1}^{*}$ (5.4)
and
$\tilde{B}^{\downarrow}_{1}\otimes\tilde{C_{1}}+\tilde{B}^{\downarrow}_{2}\otimes\tilde{C}_{2}=q_{1}q_{1}^{*}\otimes ee^{*}+2q_{1}q_{1}^{*}\otimes q_{1}q_{1}^{*}+q_{2}q_{2}^{*}\otimes q_{1}q_{1}^{*}.$ (5.5)
It can be verified that
$s_{2}(\tilde{B}^{\downarrow}_{1}\otimes\tilde{C_{1}}+\tilde{B}^{\downarrow}_{2}\otimes\tilde{C}_{2})-s_{2}(\tilde{B}_{1}\otimes\tilde{C_{1}}+\tilde{B}_{2}\otimes\tilde{C}_{2})<-0.05.$ (5.6)
The relation Eq. 5.3 holds in cases where $B_{1}$ and $B_{2}$ are both rank-$1$. This follows from a majorization relation for separable quantum states due to Nielsen and Kempe [19].
Acknowledgments
I thank Manny Knill for many hours of stimulating and encouraging discussions, and Milad Marvian for pushing me to find counterexamples. I am grateful to Felix Leditzky for comments that improved the presentation of the material. This work was supported by National Science Foundation Grant PHY-2116246.
Appendix A Remainder of the proof of Thm. 3.5
Following the strategy outlined in Sec. 3.3, we show that for every feasible pair, there exists a symmetric feasible pair that does just as well. It is useful to observe that for any $(S_{1},S_{2})\subseteq[d]\times[d]$ such that $|S_{1}|=|S_{2}|=k$, in addition to the basic constraints, certain alignment constraints are satisfied as well.
Lemma A.1.
For a $k$-set pair $(S_{1},S_{2})\subseteq[d]\times[d]$ and $(\ell_{1},\ell_{2})\in[d]\times[d]$,
$|S_{1}\cap\Omega^{1}_{\ell_{1}}|+|S_{2}\cap\Omega^{2}_{\ell_{2}}|\leq\min(2k,|\Omega^{1}_{\ell_{1}}\cap\Omega^{2}_{\ell_{2}}|+|\Omega^{1}_{\ell_{1}}\cup\Omega^{2}_{\ell_{2}}|).$ (A.1)
Proof.
The fact that $|S_{1}|=|S_{2}|=k$ implies $|\Omega^{1}_{\ell_{1}}\cap S_{1}|+|\Omega^{2}_{\ell_{2}}\cap S_{2}|\leq 2k$. The other inequality follows from inclusion-exclusion:
$|\Omega^{1}_{\ell_{1}}\cap S_{1}|+|\Omega^{2}_{\ell_{2}}\cap S_{2}|\leq|\Omega^{1}_{\ell_{1}}|+|\Omega^{2}_{\ell_{2}}|=|\Omega^{1}_{\ell_{1}}\cap\Omega^{2}_{\ell_{2}}|+|\Omega^{1}_{\ell_{1}}\cup\Omega^{2}_{\ell_{2}}|.\qed$
This means as long as we restrict to $k$-set pairs $(S_{1},S_{2})$, we need only be concerned with constraints of the form
$|S_{1}\cap\Omega^{1}_{\ell_{1}}|+|S_{2}\cap\Omega^{2}_{\ell_{2}}|\leq|\Omega^{1}_{\ell_{1}}\cap\Omega^{2}_{\ell_{2}}|+k.$ (A.2)
We start by showing that every symmetric pair of $k$-sets is feasible. This shows explicitly that the feasible set is not empty.
Lemma A.2.
If $S\subseteq[d]$ satisfies $|S|=k$, then $(S,S)$ is a feasible pair.
Proof.
Since $|S|=k$, the vector $\mathtt{1}_{S}\oplus\mathtt{1}_{S}$ satisfies the basic constraints. For $(\ell_{1},\ell_{2})\in[d]\times[d]$, using inclusion-exclusion and the fact that intersections distribute over unions yields
$|S\cap\Omega^{1}_{\ell_{1}}|+|S\cap\Omega^{2}_{\ell_{2}}|$ $=|(S\cap\Omega^{1}_{\ell_{1}})\cap(S\cap\Omega^{2}_{\ell_{2}})|+|(S\cap\Omega^{1}_{\ell_{1}})\cup(S\cap\Omega^{2}_{\ell_{2}})|$
$=|S\cap(\Omega^{1}_{\ell_{1}}\cap\Omega^{2}_{\ell_{2}})|+|S\cap(\Omega^{1}_{\ell_{1}}\cup\Omega^{2}_{\ell_{2}})|$
$\leq\min(k,|\Omega^{1}_{\ell_{1}}\cap\Omega^{2}_{\ell_{2}}|)+\min(k,|\Omega^{1}_{\ell_{1}}\cup\Omega^{2}_{\ell_{2}}|).\qed$
Let $(S_{1},S_{2})$ be an arbitrary feasible pair. We argue into two steps that it may be adjusted to a symmetric pair without decreasing the objective value. In the first step, we fix $S_{1}$ and optimize the second argument. Specifically, we find the optimal $k$-set $S\subseteq[d]$ such that $(S_{1},S)$ is a feasible pair. In the second step, we show that the objective value at $(S,S)$ is greater than or equal to the objective value at $(S_{1},S)$.
A.1 First step
Let $S_{1}=\{x_{1},\ldots,x_{k}\}\subseteq[d]$ with its elements denoted so that $x_{k}\geq^{1}\cdots\geq^{1}x_{2}\geq^{1}x_{1}$. For each $i\in[k]$, let $m_{i}\in[d]$ be minimal such that $x_{i}\in\Omega^{1}_{m_{i}}$. That means $m_{i}$ is the order of $x_{i}$ according to $\geq^{1}$. This implies $\Omega^{1}_{m_{1}}\subset\Omega^{1}_{m_{2}}\cdots\subset\Omega^{1}_{m_{k}}$.
Let $i^{\prime}\in[k]$ be given. Keeping in mind the constraints in Eq. A.2, we consider the set of all $r\in[d]$ such that for all $i\in[k]$, the constraint
$i+i^{\prime}\leq k+|\Omega^{1}_{m_{i}}\cap\Omega^{2}_{r}|$ (A.3)
is satisfied. This set is not empty because for all $i\in[k]$, the intersection $\Omega^{1}_{m_{i}}\cap\Omega^{2}_{d}={\Omega^{1}_{m_{i}}}\cap{[d]}=\Omega^{1}_{m_{i}}$, and so $k+|\Omega^{1}_{m_{i}}\cap\Omega^{2}_{d}|=k+m_{i}\geq i^{\prime}+i$. Define $r_{i^{\prime}}$ to be minimal in this set according to the $\geq^{2}$ order.
Lemma A.3.
For each $i^{\prime}\in[k]$, there exists some $i\in[k]$ such that
$i+i^{\prime}=k+|\Omega^{1}_{m_{i}}\cap\Omega^{2}_{r_{i^{\prime}}}|.$ (A.4)
Proof.
The minimality of $r_{i^{\prime}}$ implies that either $r_{i^{\prime}}=1$ or $r_{i^{\prime}}>1$ and there exists $i\in[k]$ such that $i+i^{\prime}>k+|\Omega^{1}_{m_{i}}\cap\Omega^{2}_{r_{i^{\prime}}-1}|$.
If $r_{i^{\prime}}=1$, then the constraint in Eq. A.3 corresponding to $k$ reads
$k+i^{\prime}\leq k+\underbrace{|\Omega^{1}_{m_{k}}\cap\Omega^{2}_{1}|}_{\leq 1},$ (A.5)
implying that $i^{\prime}=1$ and $|\Omega^{1}_{m_{k}}\cap\Omega^{2}_{1}|=1$.
If $r_{i^{\prime}}>1$, then since $|\Omega^{1}_{m_{i}}\cap\Omega^{2}_{r_{i^{\prime}}}|-|\Omega^{1}_{m_{i}}\cap\Omega^{2}_{r_{i^{\prime}}-1}|\leq 1$, it must hold that $i+i^{\prime}=k+|\Omega^{1}_{m_{i}}\cap\Omega^{2}_{r_{i^{\prime}}}|$. ∎
For each $i^{\prime}\in[k]$, we denote the maximal element in $\Omega_{r_{i^{\prime}}}^{2}$ according to the $\geq^{2}$ order by $y_{i^{\prime}}$. The set we seek is $S:=\{y_{1},\ldots,y_{k}\}$.
Lemma A.4.
The set $S$ is a $k$-set and its elements satisfy
$y_{k}\geq^{2}\cdots\geq^{2}y_{2}\geq^{2}y_{1}.$ (A.6)
Proof.
Let $i^{\prime}_{1},i^{\prime}_{2}\in[k]$ be such that $i^{\prime}_{1}<i^{\prime}_{2}$. By Lem. A.3, there exists at least one $i\in[k]$ such that $i+i^{\prime}_{1}=k+|\Omega^{1}_{m_{i}}\cap\Omega^{2}_{r_{i^{\prime}_{1}}}|$. Then Eq. A.3 with $i^{\prime}=i^{\prime}_{2}$ implies
$i+i^{\prime}_{1}=k+|\Omega^{1}_{m_{i}}\cap\Omega^{2}_{r_{i^{\prime}_{1}}}|<i+i^{\prime}_{2}\leq k+|\Omega^{1}_{m_{i}}\cap\Omega^{2}_{r_{i^{\prime}_{2}}}|.$
Therefore, $r_{i^{\prime}_{2}}$ has to be strictly larger than $r_{i^{\prime}_{1}}$. This shows that $|S|=k$ and $y_{k}\geq^{2}\cdots\geq^{2}y_{2}\geq^{2}y_{1}$. ∎
Lemma A.5.
The pair $(S_{1},S)$ is a feasible pair.
Proof.
By Lem. A.4, $|S|=k$ and so $(S_{1},S)$ is a $k$-set pair. Therefore, by Lem. A.1, it suffices to check that for all $(\ell_{1},\ell_{2})\in[d]\times[d]$,
$|S_{1}\cap\Omega^{1}_{\ell_{1}}|+|S\cap\Omega^{2}_{\ell_{2}}|\leq|\Omega^{1}_{\ell_{1}}\cap\Omega^{2}_{\ell_{2}}|+k.$ (A.7)
If $S_{1}\cap\Omega^{1}_{\ell_{1}}$ or $S\cap\Omega^{1}_{\ell_{2}}$ is empty, then the constraint is satisfied. Otherwise, let $i\in[k]$ be maximal such that $m_{i}\leq\ell_{1}$, and $i^{\prime}\in[k]$ be maximal such that $r_{i^{\prime}}\leq\ell_{2}$. Then
$|S_{1}\cap\Omega^{1}_{\ell_{1}}|+|S\cap\Omega^{2}_{\ell_{2}}|$ $=|S_{1}\cap\Omega^{1}_{m_{i}}|+|S\cap\Omega^{2}_{r_{i^{\prime}}}|$ (A.8)
$=i+i^{\prime}$ (A.9)
$\leq k+|\Omega^{1}_{m_{i}}\cap\Omega^{2}_{r_{i^{\prime}}}|$ (A.10)
$\leq k+|\Omega^{1}_{\ell_{1}}\cap\Omega^{2}_{\ell_{2}}|,$ (A.11)
where the last inequality follows from the containments $\Omega^{1}_{m_{i}}\subseteq\Omega^{1}_{\ell_{1}}$ and $\Omega^{2}_{r_{i^{\prime}}}\subseteq\Omega^{2}_{\ell_{2}}$. ∎
Lastly, we show that the pair $(S_{1},S)$ is better than $(S_{1},S_{2})$.
Lemma A.6.
The objective value at $(S_{1},S)$ is greater than or equal to the objective value at $(S_{1},S_{2})$.
Proof.
Let $S_{2}=\{z_{1},\ldots,z_{k}\}$, denoted so that $z_{k}\geq^{2}\cdots\geq^{2}z_{2}\geq^{2}z_{1}$. Recall that for each $i^{\prime}\in[k]$, $r_{i^{\prime}}$ is the order of the eigenvalue $\tilde{\lambda}_{y_{i^{\prime}}}(D_{2})$. Since $r_{i^{\prime}}$ is chosen to be minimal, the inequality $z_{i^{\prime}}\geq^{2}y_{i^{\prime}}$ holds. This implies
$\tilde{\lambda}_{y_{i^{\prime}}}(D_{2})-\tilde{\lambda}_{z_{i^{\prime}}}(D_{2})\geq 0.$ (A.12)
The difference between objective value at $(S_{1},S)$ and the objective value at $(S_{1},S_{2})$ is $\tilde{\lambda}(D_{2})^{T}(\mathtt{1}_{S}-\mathtt{1}_{S_{2}})$ which equals $\sum_{i^{\prime}=1}^{k}\tilde{\lambda}_{y_{i^{\prime}}}(D_{2})-\tilde{\lambda}_{z_{i^{\prime}}}(D_{2})\geq 0$. ∎
This completes the first step.
A.2 Second step
Now we show that that the objective value at $(S,S)$ is greater than or equal to the objective value at $(S_{1},S)$. We do so by showing that for each $i\in[k]$, there exist at least $i$ elements in $S$ that are smaller than or equal to $x_{i}$ in the $\geq^{1}$ order.
Lemma A.7.
For each $i\in[k]$, $|S\cap\Omega^{1}_{m_{i}}|\geq i$.
Proof.
We argue by strong induction on $p:=k-i+1$. We start with the base case $p=1$. As in the proof of Lem. A.3, for each $i^{\prime}$, there exists $i\in[k]$ such that
$i+i^{\prime}=k+|\Omega^{1}_{m_{i}}\cap\Omega^{2}_{r_{i^{\prime}}}|$ (A.13)
and either $r_{i^{\prime}}=1$, or $r_{i^{\prime}}>1$ and $|\Omega^{1}_{m_{i}}\cap\Omega^{2}_{r_{i^{\prime}}}|-|\Omega^{1}_{m_{i}}\cap\Omega^{2}_{r_{i^{\prime}}-1}|=1$. If $r_{i^{\prime}}=1$, then $i^{\prime}=1$ and $y_{1}\in\Omega^{1}_{m_{k}}$. If $r_{i^{\prime}}>1$, then $y_{i^{\prime}}\in\Omega^{1}_{m_{i}}\subseteq\Omega^{1}_{m_{k}}$. Therefore, $S\subseteq\Omega^{1}_{m_{k}}$, and so $|S\cap\Omega^{1}_{m_{k}}|=k$.
Define $\gamma(1):=k$. For $i^{\prime}\in[k]$ such that $i^{\prime}>1$, define $\gamma(i^{\prime})\in[k]$ to be minimal such that
$\gamma(i^{\prime})+i^{\prime}=k+|\Omega^{1}_{m_{\gamma(i^{\prime})}}\cap\Omega^{2}_{r_{i^{\prime}}}|>k+|\Omega^{1}_{m_{\gamma(i^{\prime})}}\cap\Omega^{2}_{r_{i^{\prime}}-1}|.$ (A.14)
Suppose that $p>1$ and the statement is true for all elements of $[p-1]$. Let $T=\{i^{\prime}_{1},\ldots,i^{\prime}_{|T|}\}$ be the subset of indices $i^{\prime}\in[k]$ that satisfy $k-\gamma(i^{\prime})+1<p$, denoted so $i^{\prime}_{1}<\ldots<i^{\prime}_{|T|}$. Notice that $i^{\prime}_{1}=1$. Observe that for $i^{\prime}\in T^{c}$, $y_{i^{\prime}}\in\Omega^{1}_{m_{\gamma(i^{\prime})}}$ implies $y_{i^{\prime}}\in\Omega^{1}_{m_{k-p+1}}$.
For each $i^{\prime}\in T$, we have the equality
$|\Omega^{1}_{m_{\gamma(i^{\prime})}}\cap\Omega^{2}_{r_{i^{\prime}}}|=\gamma(i^{\prime})+i^{\prime}-k$ (A.15)
and $k-\gamma(i^{\prime})+1<p$. By the inductive hypothesis, $|\Omega^{1}_{m_{\gamma(i^{\prime})}}\cap S|\geq\gamma(i^{\prime})$. Because of the containment $(\Omega^{1}_{m_{\gamma(i^{\prime})}}\cap S)\cap(\Omega^{2}_{r_{i^{\prime}}}\cap S)\subseteq\Omega^{1}_{m_{\gamma(i^{\prime})}}\cap\Omega^{2}_{r_{i^{\prime}}}$ and the inclusion-exclusion inequality
$|(\Omega^{1}_{m_{\gamma(i^{\prime})}}\cap S)\cap(\Omega^{2}_{r_{i^{\prime}}}\cap S)|\geq\gamma(i^{\prime})+i^{\prime}-k$ (A.16)
the equality Eq. A.15 implies
$(\Omega^{1}_{m_{\gamma(i^{\prime})}}\cap S)\cap(\Omega^{2}_{r_{i^{\prime}}}\cap S)=\Omega^{1}_{m_{\gamma(i^{\prime})}}\cap\Omega^{2}_{r_{i^{\prime}}},$ (A.17)
which implies $\Omega^{1}_{m_{\gamma(i^{\prime})}}\cap\Omega^{2}_{r_{i^{\prime}}}\subseteq S$. This in turn implies $\Omega^{1}_{m_{k-p+1}}\cap\Omega^{2}_{r_{i^{\prime}}}\subseteq S$ for all $i^{\prime}\in T$. The constraint Eq. A.2 implies
$|\Omega^{1}_{m_{k-p+1}}\cap\Omega^{2}_{r_{i^{\prime}_{|T|}}}|\geq k-p+1+i^{\prime}_{|T|}-k=i^{\prime}_{|T|}-p+1.$ (A.18)
Since the intersection $({\Omega^{1}_{m_{k-p+1}}}\cap{\Omega^{2}_{r_{i^{\prime}_{|T|}}}})$ satisfies $({\Omega^{1}_{m_{k-p+1}}}\cap{\Omega^{2}_{r_{i^{\prime}_{|T|}}}})\subseteq({\Omega^{2}_{r_{i^{\prime}_{|T|}}}}\cap S)$, it must contain at least $i^{\prime}_{|T|}-p+1+|T|-i^{\prime}_{|T|}=|T|-p+1$ elements of the set $\{y_{i^{\prime}_{1}},y_{i^{\prime}_{2}},\ldots,y_{i^{\prime}_{|T|}}\}$. Together with the elements whose subscripts are in $T^{c}$, this implies $|S\cap\Omega^{1}_{m_{k-p+1}}|\geq k-|T|+|T|-p+1=k-p+1.$
Lemma A.8.
The objective value at $(S,S)$ is greater than or equal to the objective value at $(S_{1},S)$.
Proof.
We relabel the elements of $S$ to reflect their order according to the $\geq^{1}$ order. Let us denote them with $x^{\prime}_{1},x^{\prime}_{2},\ldots,x^{\prime}_{k}$ such that $x^{\prime}_{k}\geq^{1}\cdots\geq^{1}x^{\prime}_{2}\geq^{1}x^{\prime}_{1}$.
By Lem. A.7, for each $i\in[k]$, there exist at least $i$ elements in $S$ that are in $\Omega^{1}_{m_{i}}$. In particular, this implies that $\{x^{\prime}_{1},\ldots,x^{\prime}_{i}\}\subseteq\Omega^{1}_{m_{i}}$. Hence, for all $i\in[k]$, $x_{i}\geq^{1}x^{\prime}_{i}$.
The difference between objective value at $(S,S)$ and the objective value at $(S_{1},S)$ is $\tilde{\lambda}(D_{1})^{T}(\mathtt{1}_{S}-\mathtt{1}_{S_{1}})$ which equals $\sum_{i=1}^{k}\tilde{\lambda}_{x^{\prime}_{i}}(D_{1})-\tilde{\lambda}_{x_{i}}(D_{1})\geq 0$. ∎